OpenStudy (anonymous):

Find dn/dP for (P+n^(2)*a/V^(2))(V-nb)=nRT

OpenStudy (anonymous):

which are variables and which are constants.

zepdrix (zepdrix):

Is this what your question looks like? \[\Large \frac{P+n^2}{V^2}\left(V-nb\right)=nRT, \qquad\quad\text{find}\quad \frac{dn}{dP}\] Ya, could you clarify which are constants living? I assume just \(\Large n\) and \(\Large P\) are variables maybe?

zepdrix (zepdrix):

\[\Large \frac{P+n^2a}{V^2}\left(V-nb\right)=nRT, \qquad\quad\text{find}\quad \frac{dn}{dP}\]

OpenStudy (anonymous):

\[\left( P+\frac{ n ^{2}a }{V ^{2} }\right)\left( V-nb \right)=n R T\] which are constants and which are variables.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."